# Cluster Magnification, Root Capacity, Unique Chains and Base Change

### v3

@@ 8.2. Base change and root capacity
$K^{\prime}\subset M\subset LK^{\prime}\iff$ $M=NK^{\prime}$ for a field $K\subset N\subset L$. This $N$ is unique and given by $N=\tilde{L}\cap M$.
@@
From Lemma 8.1.6, number of subgroups of ${\rm Gal}(\tilde{L}/K)$ containing ${\rm Gal}(\tilde{L}/L)=$ number of intermediate extensions for $L/K=$ number of intermediate extensions for $LK^{\prime}/K^{\prime}$. The uniqueness follows from Lemma 2.2.2. ∎
@@
Let $M/K$ be extension of $L/K$ such that $\tilde{M}$ and $K^{\prime}$ are also linearly disjoint over $K$. Then $\rho_{K}(M,L)=\rho_{K^{\prime}}(MK^{\prime},LK^{\prime})$.
@@
Let $\rho_{K}(M,L)=a_{K}\cdot r_{K}(L)$ and $\rho_{K^{\prime}}(MK^{\prime},LK^{\prime})=a_{K^{\prime}}\cdot r_{K^{\prime}}(LK^{\prime})$ where $a_{K},a_{K^{\prime}}$ are as in Lem. 6.2.5. From Lem. 8.1.6, we have $r_{K}(L)=r_{K^{\prime}}(LK^{\prime})$.
@@
Now for $L_{1}\subset\tilde{L}$, we claim $L_{1}\subset M\cap\tilde{L}\iff L_{1}K^{\prime}\subset MK^{\prime}\cap\tilde{L}K^{\prime}$. Suppose $L_{1}K^{\prime}\subset MK^{\prime}\cap\tilde{L}K^{\prime}$. Thus $L_{1}K^{\prime}\cap\tilde{M}\subset(MK^{\prime}\cap\tilde{M})\cap(\tilde{L}K^{\prime}\cap\tilde{M})$. By Lemma 2.2.2, we have $L_{1}\subset M\cap\tilde{L}$. The other implication is clear. Hence by Lem. 8.1.3, we are done.
@@ 8.3. Base change and unique chains
We study how the unique chains are related to the base change to $K^{\prime}$.
@@
$\tilde{L}^{H^{G}}=F_{1}\iff(\tilde{L}K^{\prime})^{H^{G}}=F_{1}K^{\prime}$.
@@ 9. Properties of ascending index
Let $M/K$ be obtained by strong cluster magnification with magnification factor $d$. Then $t_{K}(M)=d\ t_{K}(L)$ and $u_{K}(M)=u_{K}(L)$.
@@
Let $K$ be a number field. Let $n>2$ and $t|n$. Then there exists an extension $L/K$ of degree $n$ with ascending index $t_{K}(L)=t$.
@@
Now suppose $1<t<n$. Hence $1<n/t<n$. By Thm. 3.1.1 for $r=n/t$ and Thm. 7.3.4, there exists an $L/K$ of degree $n$ with $r_{K}(L)=n/t$ and $t_{K}(L)=t$.
@@
(2) Consider $P=\mathbb{Q}(\sqrt[4]{2})$ which has degree $4$ over $K=\mathbb{Q}$ and $t_{K}(P)=2$. For $t$ an even number, by Lemma 2 in [5], there exists $F/K$ Galois of degree $t/2$ such that $\tilde{P}$ and $F$ are linearly disjoint over $K$. Hence by Thm. 9.0.4, $L=PF$ has degree $2t$ over $K$ and $t_{K}(L)=t$.

### v4

@@ 8.2. Base change and root capacity
Collection of intermediate fields $M$ lying between $K^{\prime}$ and $LK^{\prime}$ is in bijective correspondence with collection of intermediate fields $N$ lying between $K$ and $L$ by means of the correspondence $N\mapsto NK^{\prime}$. For a given $M$, corresponding $N=\tilde{L}\cap M$.
@@
From Lemma 8.1.7, number of subgroups of ${\rm Gal}(\tilde{L}/K)$ containing ${\rm Gal}(\tilde{L}/L)=$ number of intermediate extensions for $L/K=$ number of intermediate extensions for $LK^{\prime}/K^{\prime}$. The other statement follows from Lemma 2.2.2. ∎
@@
Consider $M/L/K$. Let $M/K$ have a base change of $K$ by $K^{\prime}$. Then $\rho_{K}(M,L)=\rho_{K^{\prime}}(MK^{\prime},LK^{\prime})$.
@@
Let $\rho_{K}(M,L)=a_{K}\cdot r_{K}(L)$ and $\rho_{K^{\prime}}(MK^{\prime},LK^{\prime})=a_{K^{\prime}}\cdot r_{K^{\prime}}(LK^{\prime})$ where $a_{K},a_{K^{\prime}}$ are as in Lem. 6.2.5. From Lem. 8.1.7, we have $r_{K}(L)=r_{K^{\prime}}(LK^{\prime})$.
@@
Now for $L_{1}\subset\tilde{L}$, we claim $L_{1}\subset M\cap\tilde{L}\iff L_{1}K^{\prime}\subset MK^{\prime}\cap\tilde{L}K^{\prime}$. Suppose $L_{1}K^{\prime}\subset MK^{\prime}\cap\tilde{L}K^{\prime}$. Thus $L_{1}K^{\prime}\cap\tilde{M}\subset(MK^{\prime}\cap\tilde{M})\cap(\tilde{L}K^{\prime}\cap\tilde{M})$. By Lemma 2.2.2, we have $L_{1}\subset M\cap\tilde{L}$. The other implication is clear. Hence by Lem. 8.1.4, we are done.∎
@@ 8.3. Base change and unique chains
We study how the unique chains are influenced by base change.
@@
$\tilde{L}^{H^{G}}=F_{1}\iff(\tilde{L}K^{\prime})^{H^{G}}=F_{1}K^{\prime}$.∎
@@ 9. Properties of Ascending Index
(Ascending Index Magnification Theorem) Let $M/K$ be obtained by strong cluster magnification with magnification factor $d$. Then
$t_{K}(M)=d\ t_{K}(L)\ \text{and}\ u_{K}(M)=u_{K}(L).$
@@
(Inverse Ascending Index Problem for Number Fields) Let $K$ be a number field. Let $n>2$ and $t|n$. Then there exists an extension $L/K$ of degree $n$ with ascending index $t_{K}(L)=t$.
@@
Now suppose $1<t<n$. Hence $1<n/t<n$. By Thm. 3.1.1 for $r=n/t$ and Thm. 7.3.4, there exists an $L/K$ of degree $n$ with $r_{K}(L)=n/t$ and $t_{K}(L)=t$. ∎
@@
(2) Consider $P=\mathbb{Q}(\sqrt[4]{2})$ which has degree $4$ over $K=\mathbb{Q}$ and $t_{K}(P)=2$. For $t$ an even number, by Lemma 2 in [5], there exists $F/K$ Galois of degree $t/2$ such that $\tilde{P}$ and $F$ are linearly disjoint over $K$. Hence by Thm. 9.0.4, $L=PF$ has degree $2t$ over $K$ and $t_{K}(L)=t$. ∎